Commuter trains leaves every 20 minutes from 6 am through 9am and every 30 minutes until 9 pm. How many trains leave the station from 6 am through 4pm?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(53765)

.
3*3 + 1 = 10 trains every 20 minutes from 6 am to 9 am inclusively

    PLUS

2*7 = 14 trains every 30 minutes from 9:30 am to 4 pm inclusively.



10 + 14 = 24 trains, in total.



ANSWER.  24 commuter trains, under given conditions.

Solved.

Check if you did not forget to post your "Thanks" to me . . .


Answer by MathTherapy(10810)


Commuter trains leaves every 20 minutes from 6 am through 9am and every 30 minutes until 9 pm. How many trains leave the station from 6 am through 4pm?
Number from 6:00 a.m. - 8:40 a.m., inclusive: 9
Number from 9:00 a.m. - 4:00 p.m., inclusive: 15